Rachael has a mass of 76 kg. What is her weight in Newtons?

Rachael stands on a bathroom scale that reads in Newtons. What is the normal force exerted on Rachael by the scale?

Rachael places her hand under the sink and pulls up with a force of 72 Newtons. What is Rachael's weight now?

What is the normal force exerted on Rachael by the scale while she is pulling up the sink?
Answer #1

a] weight = mg = 76*9.8 = 744.8 N

b] Force equation on Rachel, normal - mg = 0

normal = mg = 744.8 N

c] Weight of Rachel will be same when gravity g is same, weight = mg =  744.8 N

d] Force equation on Rachel, Normal -mg - F = 0

normal = mg + F = 744.8 +72 = 816.8 N